# Learning with Noisy Labels: the Exploration of Error Bounds in Classification

### v1

@@ 5.3 Statistical error
Theorem 5.3.
Let ${\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ represent a class of the norm-constrained neural network, characterized by an input dimension $d$, an output dimension of $K$, a width and depth denoted by ${\mathcal{W}}$ and ${\mathcal{D}}$, and a constraint ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\left\{\|(W_{i},b_{i})\|_{1,\infty},1\right\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer for $i=0,1,\cdots,{\mathcal{D}}$. Let $S=\left\{z_{i}=(x_{i},y_{i})\right\}_{i=1}^{n}$ be a strictly stationary $\beta$-mixing sequence and $(a_{n},\mu_{n})$ be an integer pair such that $n=a_{n}\mu_{n}$. Suppose the loss function we consider satisfying $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$, where $\lambda>0$ is a given constant. Then we have
$\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]\leq\frac{4\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{2\lambda Kn\beta_{a_{n}}}{a_{n}}.$
@@
$\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,k}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]=\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(z_{i})-\mathbb{E}_{\tilde{S}}[\ell_{f}(\tilde{z_{i}})]\right|\right]$
$=$ $\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{S}}[\ell_{f}(z_{i})]-\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{S}}[\ell_{f}(\tilde{z_{i}})]\right|\right]$
$\leq$ $\mathbb{E}_{S,\tilde{S}}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(z_{i})-\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(\tilde{z_{i}})\right|\right]$
$=$ $2\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}\ell_{f}(\hat{z}_{i})\right|\right]$
$\leq$ $2\lambda K\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in\mathcal{F}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}f(\hat{x}_{i})\right|\right],$ (5.14)
where $\hat{S}=\{\hat{z}_{1},\cdots,\hat{z}_{2n}\}=S\cup\tilde{S}$, $\sigma_{i},i=1,\cdots,2n$ are i.i.d. Rademacher random variables that are independent of $\hat{S}$ and $\sigma=\{\sigma_{i}\}^{2n}_{i=1}$. For the last inequality, we combine Lemma 5.5 and the fact that $\ell_{f}(\hat{z}_{i})-\ell_{f}(\hat{z}^{\prime}_{i})\leq\lambda\|f(\hat{x}_{i})-f(\hat{x}^{\prime}_{i})\|_{1}\leq K\lambda|f_{j}(\hat{x}_{i})-f_{j}(\hat{x}^{\prime}_{i})|$ with $j=\mathop{\arg\max}_{j\in\{1,\cdots,K\}}|f_{j}(\hat{x}_{i})-f_{j}(\hat{x}^{\prime}_{i})|$.
In addition, we define
$\hat{x}_{1,a_{n}}$ $=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}_{i},\ \sigma_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\sigma_{i},\ y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})=\sum_{i\in G_{j}}\sigma_{i}f(\hat{x}_{i}),$
$\hat{x}_{2,a_{n}}$ $=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in H_{j}}\hat{x}_{i},\ \sigma_{2,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in H_{j}}\sigma_{i},\ y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})=\sum_{i\in H_{j}}\sigma_{i}f(\hat{x}_{i}).$
$\Xi_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}^{\prime}_{i},\ W_{j,g}(\Xi_{1,a_{n}},\sigma_{1,a_{n}})=\sum_{i\in G_{j}}\sigma_{i}f(\hat{x}^{\prime}_{i}).$
Then $\Xi_{1,a_{n}}$ is the IB sequence of $x_{1,a_{n}}$. We have
$\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}f(\hat{x}_{i})\right|\right]$
$=$ $\mathbb{E}_{\hat{S},\sigma}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})+\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})\right|\right)$
$\leq$ $\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})\right|\right)$
$\qquad+\mathbb{E}_{\hat{x}_{2,a_{n}},\sigma_{2,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})\right|\right)$
$=$ $2\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})\right|\right)$
$=$ $\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{\mu_{n}}\sum_{j=1}^{\mu_{n}}\frac{y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})}{a_{n}}\right|\right)$
$\leq$ $\mathbb{E}_{\Xi_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{\mu_{n}}\sum_{j=1}^{\mu_{n}}\frac{W_{j,f}(\Xi_{1,a_{n}},\sigma_{1,a_{n}})}{a_{n}}\right|\right)+\mu_{n}\beta_{a_{n}}$
where the second inequality to last is from Lemma 5.4, the last inequality is from Lemma 5.6 and the last equality is from $n=a_{n}\mu_{n}$. Combining (5.3) and (5.3), we conclude the result. ∎
Remark 5.4.
The condition $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$ holds for the $\ell_{1}$ loss by the triangle inequality. For the CE or the RCE loss, this condition holds when Assumption 4.1 is satisfied, as derived from Lemma 5.1. The values of $\lambda$ for different loss functions are
$\lambda=\left\{\begin{array}[]{ll}1,&\ell_{1}\text{\ loss},\\ \frac{1}{\epsilon},&\text{CE loss},\\ -A,&\text{RCE loss}.\end{array}\right.$
Theorem 5.5.
Let ${\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ represent a class of the norm-constrained neural network, characterized by an input dimension $d$, an output dimension of $K$, a width and depth denoted by ${\mathcal{W}}$ and ${\mathcal{D}}$, and a constraint ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\left\{\|(W_{i},b_{i})\|_{1,\infty},1\right\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer for $i=0,1,\cdots,{\mathcal{D}}$. We consider the noisy dataset $S^{\eta}$ as a strictly stationary $\beta$-mixing sequence with the sample size $n$ satisfying $n=a_{n}\mu_{n}$, where $(a_{n},\mu_{n})$ is an integer pair. Suppose the loss function we consider satisfying $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$, where $\lambda>0$ is a given constant. Then we have
$\mathbb{E}_{S^{\eta}}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]\leq\frac{4\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{2\lambda Kn\beta_{a_{n}}}{a_{n}}.$
@@
5.4 Approximate error
First of all, we give a definition and some lemmas, which will be useful for the proof of approximate error bound. To estimate the boundary of the approximation error, we utilize certain proofs from [SYZ22] for functions within $\mathcal{F}_{d,1}({\mathcal{W}},{\mathcal{D}})$ and extend these proofs to encompass the neural network class $\mathcal{F}_{d,K}({\mathcal{W}},{\mathcal{D}})$.
Definition 5.5.
Let $f=[f^{(1)}\ \cdots\ f^{(K)}]^{\top}$ be a map from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$, where $f^{(i)}\in C([0,1]^{d})$. Define $w_{f}(s)$ as the continuous module,
$w_{f}(s):=\left[w_{f^{(1)}}(s)\ \cdots\ w_{f^{(K)}}(s)\right]^{\top},\ \forall s\geq 0,$ (5.16)
where $w_{f^{(i)}}(s)=\left\{\sup\left|f^{(i)}(x)-f^{(i)}(y)\right|:x,y\in[0,1]^{d},\left\|x-y\right\|_{2}\leq s\right\}$. Especially, we have
$w_{f}(s\cdot j)\leq j\cdot w_{f}(s),\ \forall j\in\mathbb{N}^{+},\forall s\geq 0.$ (5.17)
Lemma 5.7 (Lemma 25 in [FJK+23]).
Let $\mathcal{F}_{d,K}({\mathcal{W}},{\mathcal{D}})$ denote the class of functions generated by a ReLU neural network with input dimension $d$, output dimension $K$, width ${\mathcal{W}}$ and depth ${\mathcal{D}}$. Suppose $g_{1}\in\mathcal{F}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1})$, $g_{2}\in\mathcal{F}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2})$, then the following properties hold:
(1) If $d_{1}=d_{2}$, $K_{1}=K_{2}$, ${\mathcal{W}}_{1}\leq{\mathcal{W}}_{2}$, ${\mathcal{D}}_{1}\leq{\mathcal{D}}_{2}$, then ${{\mathcal{F}}}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1})\subseteq{{\mathcal{F}}}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2})$.
(2) If $K_{1}=d_{2}$, then $g_{2}\circ g_{1}\in\mathcal{F}_{d_{1},K_{2}}(\max\left\{{\mathcal{W}}_{1},{\mathcal{W}}_{2}\right\},{\mathcal{D}}_{1}+{\mathcal{D}}_{2})$.
Lemma 5.8.
Under the same assumption of Lemma 5.7, define $x=(x_{1},x_{2})^{\top}$ and $g(x)=(g_{1}(x_{1}),g_{2}(x_{2}))$, then $g\in\mathcal{F}_{d_{1}+d_{2},K_{1}+K_{2}}({\mathcal{W}}_{1}+{\mathcal{W}}_{2},\max\left\{{\mathcal{D}}_{1},{\mathcal{D}}_{2}\right\})$.
Lemma 5.9 (Proposition 3.1 in [SYZ22]).
For any $N,M,d\in\mathbb{N}^{+}$, define $J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor$ and $\delta\in(0,\frac{1}{3J}]$. Then there exists a function $\psi\in\mathcal{F}_{1,1}\left(8\left\lfloor N^{1/d}\right\rfloor+3,2\left\lfloor M^{1/d}\right\rfloor+5\right)$, such that
$\psi(x)=j,\quad x\in\left[\frac{j}{J},\frac{j+1}{J}-\delta\cdot\mathds{1}_{\left\{j\leq J-2\right\}}\right],\quad j=0,1,\cdots,J-1.$
Lemma 5.10 (Proposition 3.2 in [SYZ22]).
For any $\epsilon>0$ and any $N,M\in\mathbb{N}^{+}$, define $J$ to be a positive integer satisfying $J\leq N^{2}M^{2}\left\lfloor\log_{3}(N+2)\right\rfloor$. Assume $f(x_{j})>0$ for $j=0,1,\cdots,J-1$ are samples with
$\left|f(x_{j})-f(x_{j-1})\right|\leq\epsilon,\ j\in[J].$
Then there exists $\phi\in\mathcal{F}_{1,1}(16N+30,6M+10)$ such that
[…]
Theorem 5.6.
For any continuous map $f:[0,1]^{d}\to{\mathbb{R}}^{K}$ and any $N,M\in\mathbb{N}^{+}$, there exists a function
$\phi\in\mathcal{F}_{d,K}\left(\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\},11M+18\right),$
such that $\left\|\phi\right\|_{L^{\infty}({\mathbb{R}}^{d})}\leq\left|f(\mathbf{0})\right|+w_{f}(\sqrt{d})$ and
$\left|\phi(x)-f(x)\right|\leq\left\{\begin{array}[]{ll}130\sqrt{d}w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right),&\forall x\in[0,1]^{d}\backslash\Omega([0,1]^{d},J,\delta),\\ 2w_{f}(\sqrt{d}),&\forall x\in\Omega([0,1]^{d},J,\delta),\end{array}\right.$
where $J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor$, $\forall\delta\in(0,\frac{1}{3J}]$ and
$\Omega([0,1]^{d},J,\delta)=\bigcup_{i=1}^{d}\left\{x=[x_{1},x_{2},\cdots,x_{d}]^{\top}:x_{i}\in\bigcup_{j=1}^{J}(j/J-\delta,j/J)\right\}.$

### v2

@@ Appendix A Useful Lemmas
Lemma A.4 (Lemma 25 in [FJK+23]).
Let $\mathcal{F}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ denote the class of functions generated by a ReLU neural network with input dimension $d$, output dimension $K$, width ${\mathcal{W}}$ and depth ${\mathcal{D}}$, with weights bounded by ${\mathcal{M}}$. Suppose $g_{1}\in\mathcal{F}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1},{\mathcal{M}}_{1})$, $g_{2}\in\mathcal{F}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2},{\mathcal{M}}_{2})$, then the following properties hold:
(1) If $d_{1}=d_{2}$, $K_{1}=K_{2}$, ${\mathcal{W}}_{1}\leq{\mathcal{W}}_{2}$, ${\mathcal{D}}_{1}\leq{\mathcal{D}}_{2}$, ${\mathcal{M}}_{1}\leq{\mathcal{M}}_{2}$, then ${{\mathcal{F}}}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1},{\mathcal{M}}_{1})\subseteq{{\mathcal{F}}}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2},{\mathcal{M}}_{2})$.
(2) If $K_{1}=d_{2}$, then $g_{2}\circ g_{1}\in\mathcal{F}_{d_{1},K_{2}}(\max\left\{{\mathcal{W}}_{1},{\mathcal{W}}_{2}\right\},{\mathcal{D}}_{1}+{\mathcal{D}}_{2},{\mathcal{M}}_{2}\max({\mathcal{M}}_{1},1))$.
(3) If $d_{1}=d_{2}$ and $K_{1}=K_{2}$, then $g_{1}+g_{2}\in\mathcal{F}_{d_{1},K_{1}}({\mathcal{W}}_{1}+{\mathcal{W}}_{2},\max\left\{{\mathcal{D}}_{1},{\mathcal{D}}_{2}\right\},{\mathcal{M}}_{1}+{\mathcal{M}}_{2})$.
(4) If $d_{1}=d_{2}$, define $g(x)=(g_{1}(x),g_{2}(x))$, then
$g\in{\mathcal{F}}_{d_{1}+d_{2},K_{1}+K_{2}}({\mathcal{W}}_{1}+{\mathcal{W}}_{2},\max\{{\mathcal{D}}_{1},{\mathcal{D}}_{2}\},\max\{{\mathcal{M}}_{1},{\mathcal{M}}_{2}\}).$
Lemma A.5 (Lemma A.8 in [PV18]).
Let $\beta\in(0,\infty)$, $\beta=n+\sigma$ with $n\in\mathbb{N}_{0}$ and $\sigma\in(0,1]$ , let $d\in\mathbb{N}$. For each $f\in{\mathcal{F}}_{\beta,d,B}$ and arbitrary $x_{0}\in(0,1)^{d}$, there is a polynomial $p(x)={\textstyle\sum_{\left|\alpha\leq n\right|}c_{\alpha}(x-x_{0})^{\alpha}}$ with $c_{\alpha}\in[-C\cdot B,C\cdot B]$ for all $\alpha\in\mathbb{N}_{0}^{d}$ with $\left|\alpha\right|\leq n$ and such that
$\left|f(x)-p(x)\right|\leq C\cdot B\cdot\left|x-x_{0}\right|^{\beta},\quad\text{for all $x\in[0,1]^{d}$},$
where $C=d^{n}$.
Lemma A.6 (Lemma 27 in [FJK+23]).
For any $d\geq 2$ and $k\in\mathbb{N}$, there exists $\psi\in{\mathcal{F}}_{d,1}((6K+3)d,(k+1)d,d^{7}(2d)^{k+1})$ such that:
$\left|x_{1}\cdots x_{d}-m(x)\right|\leq 3d2^{-2k},\quad x=(x_{1},\cdots,x_{d})^{\top}\in[-1,1]^{d}.$
Moreover, $m(x)=0$ if $x_{1}\cdots x_{d}=0$.
Appendix B The Lipschitz-Type Property
We start from the Lipschitz-type property for the $\ell_{p},p\geq 1$, CE and reverse CE losses.
Lemma B.1.
For the $\ell_{p},p\geq 1$ loss function, let $\tilde{a}=\mathrm{softmax}(a)$ and $\tilde{b}=\mathrm{softmax}(b)$, we have
$\left|\ell(\tilde{a},q)-\ell(\tilde{b},q)\right|\leq\sqrt{2}K\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},q\in\mathcal{Y}.$
@@ Appendix B The Lipschitz-Type Property
Let the $i$-th softmax function be $S(x)_{i}=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$, and $s=\sum_{l}e^{x_{l}}$, then we have
$\|\nabla_{x}[S(x)_{i}]\|_{2}=$ $\sqrt{\frac{(\sum_{l\neq i}e^{x_{i}}e^{x_{l}})^{2}+\sum_{l\neq i}e^{2x_{i}}e^{2x_{l}}}{s^{4}}}\leq\sqrt{2\frac{(\sum_{l\neq i}e^{x_{i}}e^{x_{l}})^{2}}{s^{4}}}$
$\leq$ $\sqrt{2\frac{(\sum_{l\neq i}e^{x_{i}}e^{x_{l}})^{2}}{(\sum_{l\neq i}e^{x_{i}}e^{x_{l}})^{2}}}=\sqrt{2},$
where the first equality used the fact that $\sum_{i}x_{i}^{2}\leq(\sum_{i}x_{i})^{2}$. Then we have
$\|\tilde{a}-\tilde{b}\|_{1}\leq\sum_{i=1}^{K}\left|\tilde{a}_{i}-\tilde{b}_{i}\right|=\sum_{i=1}^{K}\|\nabla_{x}[S(\xi)_{i}]\|_{2}\left\|a-b\right\|_{2}\leq\sqrt{2}K\|a-b\|_{2}.$
Therefore, for any $p\geq 1$, we have $|\ell(\tilde{a},q)-\ell(\tilde{b},q)|\leq\ell(\tilde{a},\tilde{b})\leq\|\tilde{a}-\tilde{b}\|_{1}\leq\sqrt{2}K\|a-b\|_{2}.$
Lemma B.2.
For the CE loss function, let $\tilde{a}=\mathrm{softmax}(a)$ and $\tilde{b}=\mathrm{softmax}(b)$, we have
$\left|\ell(\tilde{a},q)-\ell(\tilde{b},q)\right|\leq\sqrt{2}K\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},q\in\mathcal{Y}.$
@@
Let the $i$-th softmax-CE loss function be $\mathrm{CE}(x)_{i}=-\log(\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}})$, and $s=\sum_{l}e^{x_{l}}$, then we have
$\|\nabla_{x}[\mathrm{CE}(x)_{i}]\|_{2}=\sqrt{\frac{(\sum_{l\neq i}e^{x_{l}})^{2}+\sum_{l\neq i}e^{2x_{l}}}{s^{2}}}\leq\sqrt{2\frac{({\sum_{l\neq i}e^{x_{l}}}^{2})}{s^{2}}}\leq\sqrt{2\frac{{\sum_{l\neq i}e^{x_{l}}}^{2}}{{\sum_{l\neq i}e^{x_{l}}}^{2}}}=\sqrt{2},$
where the first equality used the fact that $\sum_{i}x_{i}^{2}\leq(\sum_{i}x_{i})^{2}$. For CE loss function $\ell(\tilde{a},q)=-\sum_{i=1}^{K}q_{i}\log\tilde{a}_{i}$, $\ell(\tilde{b},q)=-\sum_{i=1}^{K}q_{i}\log\tilde{b}_{i}$, there exists $\exists\xi=[\xi_{1}\cdots\xi_{K}]^{\top}$ with $\xi_{i}\in[a_{i},b_{i}],\forall 1\leq i\leq K$ such that
$\left|\ell(\tilde{a},q)-\ell(\tilde{b},q)\right|$ $=\left|\sum_{i=1}^{K}q_{i}\log\tilde{a}_{i}-\sum_{i=1}^{K}q_{i}\log\tilde{b}_{i}\right|=\left|\sum_{i=1}^{K}q_{i}[\log\tilde{a}_{i}-\log\tilde{b}_{i}]\right|$
$\leq\sum_{i=1}^{K}\left|q_{i}\right|\left|\log\tilde{a}_{i}-\log\tilde{b}_{i}\right|=\sum_{i=1}^{K}\left|q_{i}\right|\|\nabla[\mathrm{CE}(\xi)_{i}]\|_{2}\|a-b\|_{2}$
$\leq\|q\|_{2}\sqrt{2}K\|a-b\|_{2}\leq\sqrt{2}K\|a-b\|_{2}.$
Lemma B.3.
For the reverse CE loss function, let $\tilde{a}=\mathrm{softmax}(a)$ and $\tilde{b}=\mathrm{softmax}(b)$ and set $\log 0=A$  [^3] One-hot labels may cause computational problem in the reverse CE. To solve this issue, they defined $\log 0=A$. We use the same setting as the paper [WMC+19], which is the first to introduce the reverse CE. and have
$\left|\ell(\tilde{a},q)-\ell(\tilde{b},q)\right|\leq-\sqrt{2}KA\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},q\in\mathcal{Y}.$